OpenStudy (ondinana):

@pooja195 A square floor has an area of 480 square feet. Kristin wants to put molding around the entire perimeter of the floor. Approximately how much molding does she need? A. 720 feet B. 120 feet C. 90 feet D. 88 feet

OpenStudy (ondinana):

It's 120 right? :D

OpenStudy (anonymous):

idk

OpenStudy (mathmath333):



OpenStudy (mathmath333):

question is asking u the perimeter.

OpenStudy (mathmath333):

note that it is a square \(\text{area of square}=(side)^2\) \(\text{perimeter of square}=4\times side\)

OpenStudy (ondinana):

So by my answer of 120, what exactly did I find by that? o.o

OpenStudy (mathmath333):

\(\text{area of square}=(side)^2\\480=(side)^2\\side=\sqrt{480}\\ 4\times side=4\times \sqrt{480}\\ perimeter=4\times \sqrt{480}\)

OpenStudy (ondinana):

Okay, I'm gonna try and work it out now. :D

OpenStudy (mathmath333):

good luck

OpenStudy (ondinana):

I got 87.64 so since it says approximate the answer is 88 right?

OpenStudy (mathmath333):

yep.

OpenStudy (mathmath333):

so answer was \(88\) not \(120\) right ?

OpenStudy (ondinana):

Yes, thanks for your help! ^_^